Consider a clock, which when at rest produces a flash of light every second, moving away from you at (4/5)c. (a) How frequently does it flash when it is moving at (4/5)c? (b) By how much does dis- tance between you and the clock increase between flashes? c) How much longer does it take each flash to get to your eye than the previous one? (d) What, therefore, is the in- terval between the flashes vou see? You will find that the time interval between the flashes you see is much longer than merely what time-dilatiorn predicts, because successive flashes come from further and further away. This effect is known as the Doppler shift and is covered in much more detail in Chapter 7
Answer #1

lambda=wavelength

f=frequency

fnew=f(c/c+vsource)

fnew=1(c/c+4c/5)

=5/9 hz

Now it flashes after every 9/5s

Given initial frequency(f)=1hz

vsource=4c/5

c=f(lambda)

lambda=c

lambdanew=(lambda)c/c-v

lambdanew=c^2/c-v

=5c

delta lamda=5c-c=4c